In a Young's double slitexperiment m(th) order and n(th)order of bright fringes are formed at point P on a distant screen if monochromatic source of wavelength 400nm and 600nm are used respectively.The minimum value of m and n are respectively
1
Expert's answer
2019-12-27T08:57:26-0500

Fringe width is given by "\dfrac{D}{d}\lambda"

Position of nth bright fringe is given by "\dfrac{nD}{d}\lambda"1 .......(1)

Position of mth bright fringe is given by "\dfrac{mD}{d}\lambda"2 ........(2)

then it is given that both mth and nth are formed at point P

hence, equation (1)=(2)

"\implies" "\dfrac{nD}{d}\lambda_1= \dfrac{mD}{d}\lambda_2"

"\implies" "n{\lambda}_1=m\lambda_2"

"\implies" "600n=400m"

"\implies" "3n=2m"

"{\implies} n= \dfrac{2}{3}m"

hence, putting m=3 we will get n=2

So, the minimum values of n and m are 2 and 3 respectively.